OpenStudy (anonymous):

You have 297mL of a 1.25M potassium solution, but you need to make a 0.50M potassium chloride solution. How many milliliters of water must you add to the original 1.25M solution to make a 0.50M potassium Chloride solution.

OpenStudy (anonymous):

First you have to understand what molarity is .. Molarity = amount of solute(in moles,n) / volume of solution(in liters) 1.25 M = n / 0.297L n = 0.297*1.25 = 0.371 mole if you add water into solution it's volume increases..because we don't know how many liters of water we should add...we consider V(H2O) = X liters. final molarity is 0.5M 0.5M = 0.371mole/(0.297 L + X L) by solving this eq we find x = 0.445 L or 445 mL of water

OpenStudy (anonymous):

Thanks for the help, but I don't understand the last equation. Do you divide 0.5 by .371 then multiply by .297? I couldn't get .445 when I did it. Thanks again!

OpenStudy (anonymous):